Circular No. 7957

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2001kf, 2001kg, 2001kh R. Ellis, California Institute of Technology; R. Pain, J. Raux, G. Sain-

ton, and K. Schahmaneche, Laboratoire de Physique Nucléaire et de Hautes Energies, Universities of Paris 6 and 7; and M. Sullivan, University of Durham, report the discovery of three supernovae found on images from 2001 Oct. 10 with the CFH12k camera at the Canada-France-Hawaii Telescope (CFHT) and confirmed four nights later with the CFHT. Spectral identification was obtained with a Keck telescope (+ ESI) on Oct. 20 and 23, with additional spectroscopy and imaging obtained at CFHT on Nov. 12 and 16 and at Keck on Nov. 16.

SN 2001 UT α_{2000} δ_{2000} R z Type
2001kf Oct. 10 22^{h}21^{m}34.^{s}20 + 0^{o}16′02′′.2 23.1 0.057 II-P
2001kg Oct. 10 22 20 52.51 + 0 43 05.8 22.7 0.280 Ia
2001kh Oct. 10 2 25 15.56 − 5 23 20.0 21.0 0.220 Ia

COMET 57P/DU TOIT-NEUJMIN-DELPORTE Further to his report on IAUC 7946, Z. Sekanina writes: “Using the

code developed by P. W. Chodas and myself, I was able to link the observations of the Aug. 7 secondary nucleus (MPEC 2002-P75) with those of companion F on July 17 (MPEC 2002-P30). An excellent fit (mean residual 0′′.26) suggests that F separated from nucleus A most probably in the second half of May 2001, a little more than 400 days before perihelion at a heliocentric distance of ∼ 3.6 AU. By contrast, all attempts to fit the offsets on the assumption of a separation near the 1996 perihelion have failed, leaving systematic residuals of a few arcsec. The 2001 solution is rather insensitive to the adopted nongravitational deceleration, if it is on the order of several units of 10^{−4} solar attraction. The derived separation velocity is then ∼ 5 m/s, mostly in the orbital plane. The predicted separation distances and position angles (0^{h} TT): 2002 Aug. 14, 401′′, 259^{o}.0; 24, 398′′, 259^{o}.0; Sept. 3, 393′′, 259^{o}.0; 13, 387′′, 258^{o}.8; 23, 383′′, 258^{o}.4; Oct. 3, 381′′, 257^{o}.8. Confirmation observations are needed to understand the fragmentation sequence, because major deviations from this ephemeris would indicate a different splitting scenario.”

Total visual magnitude estimates: July 28.87 UT, 12.9 (M. Reszelski, Szamotuly, Poland, 0.41-m reflector); Aug. 8.88, 13.9 (W. Hasubick, Buchloe, Germany, 0.44-m reflector); 12.94, 13.3 (R. J. Bouma, Groningen, The Netherlands, 0.31-m reflector).

2002 August 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green